a student majoring in accounting is trying to decide on the number of firms to which he should apply. given his work experience and grades he can expect to receive a job offer from 70% of the firms to which he applies. The student decides to apply to only four firms. what is the probability that he receives no job offers? carry to the nearest ten thousandths.
Answer #1

This is a direct application of Binomial distribution.

Let X be a number of job offerd receives among 4 firms

Here, X ~ Binomial ( n = 4, p = 0.70)

Probability mass function of X is,

P(X = x) = nCx px (1-p)n-x

We want to find, P(X = 0)

P(X = 0) = 4C0 (0.70)0 (1-0.70)4-0 = 1 * 1 * (0.30)4 = 0.0081

P(X = 0) = 0.0081

Hence, the probability that he receives no job offers is 0.0081